OpenStudy (anonymous):

Using the equations and enthalpy values provided, which mathematical expression can be used to determine the unknown enthalpy change represented by x? 3 C + 3 O2 yields 3 CO2 deltaH = -1182 kJ 4 H2 + 2 O2 yields 4 H2O deltaH =-1144 kJ C3H8 + 5 O2 yields 3 CO2 + 4 H2O deltaH = x C3H8 yields 3 C + 4 H2 deltaH = 104 kJ

OpenStudy (anonymous):

A)104 = (-1182) - [x + (-1144)] B)104 = x - [(-1182) + (-1144)] C)104 = (-1182) + (-1144) + x D) 104 = x + (-1182) - (-1144)